# Optimal Procurement Design: A Reduced Form Approach

### v1

@@ C.2 Proof of Theorem 2
To show that bidding the quality $q$ is a weakly dominant strategy for a seller with quality $q\in(\underline{b}_{i},\bar{b}_{i})$ for some $i=1,\ldots,M$, it suffices to compare bidding $q$ with bidding $B$, as all other possible bids are covered by the proof of Lemma 3. If she bids her quality $q$, as described in the proof of Lemma 3, her payoff is at least zero when $b^{\min}_{-}\leq\bar{b}_{M}$. When $b^{\min}_{-}=B$ and $k$ other sellers bid $B$, with probability $1-\zeta$ all other sellers are disqualified (i.e., no other sellers submit a legit bid), and therefore this seller gets $b^{\min}_{-}-q=B-q$; with probability $\zeta$, this seller’s payoff is given by $\zeta\left(\frac{B+k\bar{b}_{M}}{k+1}-q\right)$. If she bids $B$ instead, with probability $1-\zeta$ she is disqualified and gets zero payoff, and with probability $\zeta$ her payoff is $\zeta(B-q)/(k+1)$, which is strictly worse than bidding $q$ either way. This shows that bidding the quality $q$ is a weakly dominant strategy for a seller with quality $q\in(\underline{b}_{i},\bar{b}_{i})$ for some $i=1,\ldots,M$. An analogous argument shows that for a seller with quality $q\in[\bar{b}_{i-1},\underline{b}_{i}]$ for some $i=1,\ldots,M$, a weakly dominant strategy is bidding $\underline{b}_{i}$. This completes the proof. ∎
Proof of Theorem 2.
@@
Using Lemma 4, the proof of Theorem 1 indicates that the interim allocation probability induced by the aBRA, $P_{aBRA}$, equals to $\hat{P}$ defined by (11) in Proposition 2 $F$-almost everywhere on $[0,1]\setminus[S_{+},S_{0})$. For a seller with quality quantile $s\in[S_{+},S_{0})$, Lemma 4 implies that it is a weakly dominant strategy for her to bid $B$. With probability $1-\zeta$, this seller is disqualified, which means that she wins with probability 00. With probability $\zeta$, this seller is qualified; following the same steps as in the proof of Theorem 1, we see that she wins the auction with probability $\int_{S_{+}}^{S_{0}}(1-s)^{n-1}\,\mathrm{d}s/(S_{0}-S_{+})$. Thus, the interim allocation probability for this seller is
@@
C.3 Proof of Corollary 2
We first claim that $-F/f$ is decreasing. To see this, differentiate to obtain $(-F/f)^{\prime}=-1-(-Ff^{\prime}/f^{2})$. When $f^{\prime}(q)\leq 0$, because $1-F$ is log-concave, $-Ff^{\prime}/f^{2}\geq(1-F)f^{\prime}/f^{2}\geq-1$; if instead $f^{\prime}(q)>0$, $-Ff^{\prime}/f^{2}\geq-1$ since $F$ is log-concave. Thus, $(-F/f)^{\prime}\leq 0$. Then since $v(q)-q$ is strictly decreasing, for any $\gamma\in[0,1]$, both $h_{\gamma}(q)$ and $g(q)$ are strictly decreasing, and so are $\widetilde{g}$ and $\widetilde{h}_{\gamma}$. Therefore, the virtual surplus of the Lagrangian, $\widetilde{\phi}_{\gamma}(s;\lambda)=\widetilde{h}_{\gamma}(s)+\lambda\widetilde{g}(s)$ must be strictly decreasing for any $\lambda\geq 0$ and $\gamma\in[0,1]$. Consequently, for any $\lambda\geq 0$, there must exist a unique $\bar{S}\in[0,1]$ such that $\widetilde{\phi}(\bar{S};\lambda)=0$. The result thus follows.
C.4 Proof of Corollary 3
@@
Now suppose $\overline{\phi}_{\gamma}(0;\lambda^{*})=m\geq 0$. Because both $v(q)-q$ and $-F(q)/f(q)$ are concave under our assumptions, so is $\phi_{\gamma}(q;\lambda^{*})$ since it is a nonnegative linear combination of $v(q)-q$ and $-F(q)/f(q)$. Thus, $\phi_{\gamma}(q;\lambda)$ is single-peaked. Consequently, there exists $c\in[0,1]$ such that $\overline{\phi}_{\gamma}(\,\cdot\,;\lambda^{*})=m$ on $[0,c]$ and decreasing on $[c,1]$. Then by Theorem 2, if $S_{+}=S_{0}$, a BRA with one bid interval $[\underline{b},S_{0}]$ is optimal; otherwise, an aBRA with one standard bid interval $[\underline{b},F^{-1}(S_{+})]$ and extra bid $F^{-1}(S_{0})$ is optimal.

### v2

@@ C.2 Implementation: Augmenting BRA if necessary
To show that bidding the quality $q$ is a weakly dominant strategy for a seller with quality $q\in(\underline{b}_{i},\overline{b}_{i})$ for some $i=1,\ldots,M$, it suffices to compare bidding $q$ with bidding $B$, as all other possible bids are covered by the proof of Lemma 3. If she bids her quality $q$, as described in the proof of Lemma 3, her payoff is at least zero when $b^{\min}_{-}\leq\overline{b}_{M}$. When $b^{\min}_{-}=B$ and $k$ other sellers bid $B$, with probability $1-\zeta$ all other sellers are disqualified (i.e., no other sellers submit a valid bid), and therefore this seller gets $b^{\min}_{-}-q=B-q$; with probability $\zeta$, this seller’s payoff is given by $\zeta\left(\frac{B+k\overline{b}_{M}}{k+1}-q\right)$. If she bids $B$ instead, with probability $1-\zeta$ she is disqualified and gets zero payoff, and with probability $\zeta$ her payoff is $\zeta(B-q)/(k+1)$, which is strictly worse than bidding $q$ either way. This shows that bidding the quality $q$ is a weakly dominant strategy for a seller with quality $q\in(\underline{b}_{i},\overline{b}_{i})$ for some $i=1,\ldots,M$. An analogous argument shows that for a seller with quality $q\in[\overline{b}_{i-1},\underline{b}_{i}]$ for some $i=1,\ldots,M$, a weakly dominant strategy is bidding $\underline{b}_{i}$. ∎
Theorem 3 establishes the optimality of aBRA.
Theorem 3.
Let $\lambda^{*}$ be the Lagrangian multiplier associated with the optimal interim allocation identified in Proposition 3. If $\phi_{\gamma}(q;\lambda^{*})=h_{\gamma}(q)+\lambda^{*}g(q)$ is structured, then
  1. (a)
  2. (b)
Proof.
@@
Part (a) follows from Theorem 1, and it suffices to prove Part (b). Using Lemma 4, the proof of Theorem 1 indicates that the interim allocation probability induced by the aBRA, $P_{aBRA}$, equals $\hat{P}$ defined by (11) in Proposition 3 $F$-almost everywhere on $0,1]\setminus[S_{+},S_{0})$. For a seller with quality quantile $s\in[S_{+},S_{0})$, Lemma 4 implies that it is a weakly dominant strategy for her to bid $B$. With probability $1-\zeta$, this seller is disqualified, which means that she wins with probability 0. With probability $\zeta$, this seller is qualified; following the same steps as in the proof of Theorem 1, we see that she wins the auction with probability $\int_{S_{+}}^{S_{0}}(1-s)^{n-1}\,\mathrm{d}s/(S_{0}-S_{+})$. Thus, the interim allocation probability for this seller is
@@
Corollary 3.
Suppose v(q)qv(q)-q is strictly decreasing, $F$ is twice continuously differentiable, and both $F$ and 1F1-F are log-concave. Then a second-price auction with reserve price $F^{-1}(S_{0})$ is optimal.
Proof.
We first claim that F/f-F/f is decreasing. To see this, differentiate to obtain $(-F/f)^{\prime}=-1-(-Ff^{\prime}/f^{2})$. When $f^{\prime}(q)\leq 0$, because 1F1-F is log-concave, $Ff^{\prime}/f^{2}\geq(1-F)f^{\prime}/f^{2}\geq-1$; if instead $f^{\prime}(q)>0$, $Ff^{\prime}/f^{2}\geq-1$ since $F$ is log-concave. Thus, $(-F/f)^{\prime}\leq 0$. Then since v(q)qv(q)-q is strictly decreasing, for any $\gamma\in[0,1]$, both $h_{\gamma}(q)$ and $g(q)$ are strictly decreasing, and so are $\widetilde{g}$ and $\widetilde{h}_{\gamma}$. Therefore, the virtual surplus of the Lagrangian, $\widetilde{\phi}_{\gamma}(s;\lambda)=\widetilde{h}_{\gamma}(s)+\lambda\widetilde{g}(s)$ must be strictly decreasing for any $\lambda\geq 0$ and $\gamma\in[0,1]$. Consequently, for any $\lambda\geq 0$, there must exist a unique $S_{0}\in[0,1]$ such that $\widetilde{\phi}(S_{0};\lambda)=0$. The result thus follows. ∎
Corollary 3 specifies that if the buyer values marginal quality uniformly less than the potential sellers do, under standard distributional assumptions in the mechanism design literature, a second-price auction (with a reserve price if needed) maximizes any weighted average of the buyer’s expected payoff and the social surplus.
Corollary 4.
If $v(q)$ is concave and $F/f$ is convex in $q$, then $\phi_{\gamma}(q;\lambda^{*})$ is single-peaked for all $\lambda^{*}\geq 0$. Consequently,
  • if there do not exist $0\leq S_{+}<S_{0}\leq 1$ such that $\phi_{\gamma}(\cdot;\lambda^{*})=0$ on $S_{+},S_{0}]$, the BRA with one bid interval is optimal;
  • otherwise, an aBRA with one standard bid interval $\underline{b},F^{-1}(S_{+})]$ and extra bid $F^{-1}(S_{0})$ is optimal.
Proof.
@@
Now suppose $\overline{\phi}_{\gamma}(0;\lambda^{*})=m\geq 0$. Because both v(q)qv(q)-q and F(q)/f(q)-F(q)/f(q) are concave under our assumptions, so is $\phi_{\gamma}(q;\lambda^{*})$ since it is a non-negative linear combination of v(q)qv(q)-q and F(q)/f(q)-F(q)/f(q). Thus, $\phi_{\gamma}(q;\lambda)$ is single-peaked. Consequently, there exists $c\in[0,1]$ such that $\overline{\phi}_{\gamma}(\,\cdot\,;\lambda^{*})=m$ on $[0,c]$ and decreasing on $[c,1]$. Then by Theorem 3, if $S_{+}=S_{0}$, a BRA with one bid interval $\underline{b},F^{-1}(S_{0})]$ is optimal; otherwise, an aBRA with one standard bid interval $\underline{b},F^{-1}(S_{+})]$ and extra bid $F^{-1}(S_{0})$ is optimal. ∎
Assuming that $v$ is concave implies that the buyer’s marginal valuation of quality is decreasing. As pointed out by lpv, the convexity of $F/f$ is satisfied by many familiar distributions with bounded supports, including power distributions, [^17] The CDF of a power distribution takes the form of $F(x)=x^{\alpha}$, where $\alpha>0$. When $\alpha=1$, we get the uniform distribution. (truncated) Pareto distributions, and (truncated) exponential distributions. In fact, it is also satisfied by Beta distributions with both parameters greater than or equal to 1.
Appendix D An extension of the main model
In the main model, the buyer’s valuation is assumed to be a deterministic function of the sellers’ quality. Moreover, a seller’s cost, or her reservation value, is identical to her quality. These assumptions are made to simplify notation and emphasize the quality concerns. In many relevant applications, however, it might be more natural either to assume that the buyer’s valuation is a random variable or to treat the sellers’ costs as primitives, or both. In what follows, we show that, even in these settings, focusing on the main model is without loss. [^18] For brevity, we only discuss the buyer-optimal problem here; extending the analysis to allow for any weighted average is straightforward.
For concreteness, consider a buyer who would like to contract with one of several potential suppliers to develop a new project, say a new production line. The cost of supplier $s$, $c_{s}\in[0,1]$, is her private information; the costs are independently and identically distributed according to a continuous density function $f_{C}(\cdot)$. The project’s value is not perfectly revealed to the buyer until the end of the development phase at the earliest, which is long after penning the contract. Consequently, at the time of contracting the buyer’s valuation is a random variable $\Xi$. We assume that the realization of $\Xi$, $\xi\in[\underline{\xi},\overline{\xi}]$, is not contractible. The buyer believes that $\Xi$ and $C$ are correlated, and that the conditional distribution of $\Xi$ is $f_{\Xi\,|C}(\cdot|c)$. One possible assumption can be that $\Xi$ and $C$ are positively affiliated, or equivalently $\text{MTP}_{2}$ (kr80a, kr80a; see also mw82, mw82).
Let $\bm{c}\coloneqq(c_{1},\ldots,c_{n})$. Given a direct mechanism $\{p_{s}(\bm{c}),t_{s}(\bm{c})\}_{s=1}^{n}$, where for each cost profile $\bm{c}$, $p_{s}(\bm{c})$ specifies the probability that the buyer contracts with supplier $s$, and $t_{s}(\bm{c})$ is the transfer that the buyer pays to supplier $s$, the buyer’s expected payoff can be written as
where $f^{n}(\bm{c}):=\prod_{s=1}^{n}f_{C}(c_{s})$. If we define $v(c_{s}):=\mathbb{E}[\Xi\,|\,C=c_{s}]$, we see from (1) that the problem here is identical to the procurement problem we study in the main text, and the curvature of $v(c_{s})$ is governed by the conditional distribution. For example, if $\Xi$ and $C$ are positively affiliated, $v(\cdot)$ is increasing.
Example 3.
A manufacturer would like to procure a machine for production. For simplicity, suppose that her valuation is identical to the durability of the machine. She believes that the two potential sellers’ costs are identically, independently, and uniformly distributed on $[0,1]$. Conditional on the cost realization $c$, her valuation $\Xi$ is distributed according to a Pareto distribution with scale 0.5 and shape 2.2c2.2-c. [^19] The scale parameter can be interpreted as the length of the machine’s warranty. Consequently,
$v(c)=\mathbb{E}[\Xi\,|\,C=c]=\frac{1.1-0.5c}{1.2-c},$
and g(c)=v(c)2cg(c)=v(c)-2c. By Proposition 1, the optimal interim allocation is given by
$\hat{P}(c)=\left\{\begin{array}[]{ll}1-c&c<0.48,\\ 0.26&c\geq 0.48.\end{array}\right$.
By Corollary 1, a BRA with two bid intervals $[0,0.48]$ and $\{1\}$ is optimal. $\lozenge$